Sove for x, write answer as an inequality.
(2x+5)/(x-7) Can someone tell me the steps to do this problem?Answer is less than or equal to 0

Answer by uma(370)

(2x+5)/(x-7) < or = 0
multiplying both the sides by (x-7) we have,
(2x+5) < or = 0
Adding -5 to both the sides of the inequality,
2x + 5 - 5 < or = -5
==> 2x < or = -5
==>2x/2 < or = -5/2 [dividing by 2 throughout]
==> x < or = -5/2
Good Luck!!!